OpenStudy (anonymous):

What is the equation of the parabola with a directrix at y = 4 and focus (0, -4)

OpenStudy (callisto):

Just try..... ''The standard form is (x - h)^2 = 4p (y - k), where the focus is (h, k + p) and the directrix is y = k - p.'' Given: focus = (0, -4); directrix is at y=4 h=0 k+p = -4 -(1) k-p = 4 -(2) (1) + (2) 2k = 0 k =0 Put k=0 into (1) 0 + p = -4 p =-4 So, (x - h)^2 = 4p (y - k) (x-0)^2 = 4 ( -4) (y-0) x^2 = -16y I'm not sure if it is correct since I haven't learnt it before.

OpenStudy (amistre64):

looks fine to me .... can always use the wolf to dbl chk tho